OpenStudy (anonymous):

You have 3/4 of a pie leftover after your party. Each guest takes home 1/6 of the leftover pie. What fraction of the pie does each guest take home?

OpenStudy (ahsome):

Again, multiply the fractions \[\frac{3}{4}\times \frac{1}{6}\]\[\frac{3}{24}\]\[\frac{1}{8}\]